OpenStudy (anonymous):

why does sin(x) = 2 have no solutions?

OpenStudy (blockcolder):

What is the maximum value of sin(x)?

OpenStudy (anonymous):

i dont know

OpenStudy (blockcolder):

The maximum value of sin(x) is 1 and thus sin(x) can never reach 2. :D

OpenStudy (anonymous):

hmm~ how can you say the maximum value of sin(x) is 1? any proof?

OpenStudy (blockcolder):

Well, if you have some knowledge in calculus, I could explain. :D

OpenStudy (blockcolder):

Or we can look at the unit circle.

OpenStudy (anonymous):

\[\sin\theta = \frac ph\]